OpenStudy (anonymous):

f(x)=5x+2 to evaluate f(x)-f(5) I'm not sure if its -20x+4 or just -20x.

OpenStudy (anonymous):

f(x)=5x+2 f(5)= 25 +2 = 27 => f(x) - 27 = 5x - 25